# VeOmni: Scaling Any Modality Model Training with Model-Centric Distributed Recipe Zoo

### v2

@@ 1 Introduction
Extending existing training frameworks to omni-modal LLMs is non-trivial. To address the growing parameter sizes, modern training frameworks leverage various distributed training strategies [51, 40, 32, 41] to alleviate computation and memory bottlenecks. For example, Megatron-LM [51, 41] provides optimized transformer blocks [61] with advanced parallelization strategies like tensor parallelism (TP) and pipeline parallelism ($P$). However, as omni-modal architectures become complex in both functionality and parameter size, directly applying these techniques to omni-modal LLMs often leads to load imbalance and poor scalability [17], due to the current frameworks tightly couple model definition with parallel logic. This entanglement hiders generalization to more diverse model architectures. Although recent efforts [24, 77, 17] attempt to mitigate the inefficiencies of omni-modal training, they still suffer from the coupling of communication and computation, resulting in significant engineering cost and limited extensibility.
To efficiently train large omni-modal LLMs, VeOmni introduces model-centric distributed recipes that decouple model definition from parallel logic. As shown in Figure 1, distributed strategies such as fully sharded data parallel (FSDP), hybrid sharded data parallel (HSDP) [78, 36], sequence parallelism (SP) [29], and expert parallelism (EP) [74] can be applied to model blocks via a high-level parallel plan API. This design effectively separates communication from computation and has been validated in recent LLM training systems [36]. VeOmni supports flexible composition of parallel strategies (e.g., FSDP+SP for 2D and FSDP+SP+EP for 3D parallelism), enabling a range of training recipes tailored to dense or MoE models. By decoupling parallel logic, new modality-specific modules can be integrated with minimal engineering effort, as computation modules need not account for distributed concerns. In addition, VeOmni provides a lightweight interface for customizing omni-modal LLMs, allowing users to easily add or remove modality-specific encoders and decoders.
@@
(1) We propose model-centric distributed recipes that efficiently scale omni-modal training using n-D parallelism with minimal engineering overhead.
(2) We introduce a light-weight configuration interface for easy customization of omni-modal LLMs.
(3) We demonstrate VeOmni’s competitive efficiency and scalability across 8–128 GPUs on models ranging from 7B to 72B parameters under omni-modal scenarios.
@@ 2.1 Multi-Modal and Omni-Modal LLMs
Recent advances in large language models (LLMs) have increasingly focused on developing multi-modal and omni-modal capabilities. The prevailing approaches include incorporating modality-specific encoders into the input space of the LLMs for multimodal understanding [39, 65, 72, 37, 19], and attaching generative decoders to the output space for multimodal generation [27, 59] or embodied action prediction [2, 31]. More recently, omni-modal LLMs that support unified understanding and generation across modalities have emerged, aiming to align arbitrary modality features with language in a shared latent space. Based on a unified paradigm of auto-regressive modeling over multimodal embeddings, these models differ significantly in how the modality features are encoded and decoded. Some [56, 67] adopt discrete-token generation pipelines based on VQ-VAE series [16, 46]. Some [64, 25, 54] employ continuous-token generation via latent diffusion models [48, 44]. Others [12, 71, 80] explore hybrid architectures that combine diffusion-based generation with auto-regressive decoding. In addition, some works proposed alternative next-token prediction schemes, such as masked generation [34], next patch prediction [42], next scale prediction [58], next block prediction [47], offering increased flexibility beyond standard auto-regressive decoding. Given this diverse landscape of model architectures, VeOmni offers a flexible and extensible framework for building and scaling new modeling paradigms.
@@ 2.2 LLM Training Frameworks
The landscape of large language model training frameworks has evolved significantly to address the computational demands of scaling. For pure text training, several mature frameworks have established themselves as industry standards. Megatron-LM [51, 41] pioneered optimized transformer blocks with advanced parallelization strategies including tensor parallelism (TP) and pipeline parallelism ($P$), becoming the foundation for many subsequent frameworks. Colossal-AI [33] extends this paradigm by offering comprehensive 3D parallel training strategies that combine data, tensor, and pipeline parallelism for enhanced scalability. NeMo [50] provides an end-to-end cloud solution specifically designed for training large-scale LLMs with enterprise-grade features, while TorchTitan [36] and veScale [62] represent the newer generation of PyTorch-native frameworks that emphasize auto-parallelization and simplified programming models.
@@
3 VeOmni: Scalable Omni-Modal Training
@@
3.1 Light-Weight Omni-Modal Customization
Figure 2: Composite Architecture for Omni-Modal LLMs. The architecture consists of three fully decoupled modules: Encoder, Foundation, and Decoder.
To support training across arbitrary modalities, we propose a plug-and-play architecture that allows any combination of multimodal encoders and decoders to be flexibly attached to the input and output sides of a foundation model, as shown in Fig. 2. This modular and fully decoupled design decouples system-level operations from model-specific computation, allowing the streaming pipeline to operate independently of the model pipeline. Therefore, VeOmni enables easy extension and scaling of diverse omni-modal LLMs.
@@ 3.1 Light-Weight Omni-Modal Customization
During training, for input modalities, each encoder implements an lm_encode function that encodes raw modality data into token embeddings, which are then inserted into the input embeddings of the foundation model. Similarly, decoder modules implement the lm_encode function to provide training-time inputs representing the target output tokens. The final outputs from the foundation model are decoded into the target modality using lm_head function, completing the end-to-end mapping.
In the inference phase, for each prediction step, the output hidden states is passed through the corresponding modality decoder’s lm_embed function, which generates embeddings that are used as inputs for the foundation model’s next token prediction. Once all tokens have been predicted, the decoder’s lm_generate function is invoked to generate the final modality-specific output.
The details of the training and inference processes are listed in Appendix B.3.
@@ 3.2 Model-Centric Distributed Recipe Zoo
To support large-scale distributed training of omni-modal LLMs, VeOmni offers a modular suite of distributed training strategies, including Fully Sharded Data Parallelism (FSDP), Sequence Parallelism (SP), and Expert Parallelism (EP), designed to handle challenges such as training with ultra-long sequences and efficiently scaling large mixture-of-experts (MoE) models. VeOmni enables users to freely compose distributed training recipes tailored to their model architectures and modality-specific requirements, without needing to manage low-level implementation details. The framework is designed with the following principles:
@@
    Non-intrusive distributed training APIs. VeOmni decouples the model architecture from the underlying distributed training mechanisms, enabling users to compose multimodal models flexibly without modifying low-level parallelization code.
@@
    Support for long context training. VeOmni accommodates training workloads involving extremely long sequences across different modalities, which is a common challenge in omni-modal training.
@@
3.2.1 Fully and Hybrid Sharded Data Parallel
@@
Fully Sharded Data Parallel (FSDP) is a highly efficient implementation of the Zero Redundancy Optimizer (ZeRO-3) [45] in PyTorch. Its primary advantage is the significant reduction in memory required on each GPU during training. By sharding a model’s parameters, gradients, and optimizer states across all available devices, FSDP allows for the training of extremely large models that would otherwise exceed the memory capacity of a single GPU. A key advantage of FSDP is its non-intrusive design, which decouples the model’s architecture from the parallelization strategy. This means developers can focus on designing their models without needing to modify the underlying code to accommodate the distributed training setup. This non-intrusive nature makes FSDP exceptionally well-suited for training omni-modal LLMs. Since these models are defined by their complex and heterogeneous architectures, FSDP’s ability to parallelize training without requiring any changes to the model’s code makes it an ideal solution. VeOmni integrates both FSDP1 [78] and FSDP2 [36] as fundamental components of its distributed training stack, and provides a unified API for easy configuration. More details in Appendix B.5.
@@
To further improve training efficiency, VeOmni integrates Hybrid Sharded Data Parallel (HSDP), an extension of FSDP designed to minimize communication overhead. HSDP utilizes a 2D device mesh, combining FSDP within “shard groups” and Distributed Data Parallel (DDP) across “replicate groups”. This hybrid approach drastically cuts down on inter-node communication, enabling even greater scalability. The switch to the more efficient HSDP is as simple as changing one parameter in the configuration, with no modifications to the underlying code required.
3.2.2 Sequence Parallelism for Long Context Training
@@
With the advancement of state-of-the-art omni-modal LLMs, the sequence lengths required for training have grown significantly, particularly in domains such as high-resolution image or video understanding and generation [65, 76, 18]. This trend toward longer sequences poses substantial challenges in terms of both computational cost and memory consumption. Efficiently addressing the demands of long context training is therefore critical for scaling model capacity and unleashing the full potential of omni-modal architectures. To address this, VeOmni adopts DeepSpeed Ulysses [29], a highly efficient sequence parallelism technique specifically designed for transformer training on ultra-long sequences. DeepSpeed Ulysses splits activations along the sequence dimension and strategically orchestrates all-to-all communications during attention computation, ensuring that communication volume remains constant when both sequence length and device count are scaled proportionally. This design enables high efficiency and scalability for long-context model training. A core advantage of VeOmni is its provision of a non-intrusive interface for deploying Ulysses. Developers can seamlessly leverage DeepSpeed Ulysses without needing to modify model-building code or directly interact with the underlying implementation details. In particular, VeOmni enhances the implementation of Flash Attention [9] by integrating DeepSpeed Ulysses while keeping fully compatible with the default attention, enabling the straightforward adoption of advanced attention acceleration methods within sequence parallel workflows. More details in Appendix B.6.
@@ 3.2.3 Expert Parallelism for MoE Model Scaling
3.2.4 n-D Parallelism Training Strategies
Figure 3: Overview of VeOmni’s n-D Parallelism Design. The figure on the left illustrates the data flow of VeOmni during the training of omni-modal LLMs. The encoder for each modality processes its respective input, and then scatters the features to the corresponding ranks through an all-to-all communication operation. The right figure shows the 3D parallelism system of VeOmni. Attention blocks apply HSDP with data sharded size 2 and data replicate size 2, input apply Ulysses with sequence parallel size 2, Mixture-of-Experts blocks apply expert parallel and FSDP with experts parallel size 2 and data sharded size 2.
In VeOmni, core parallelism strategies (i.e., Fully Sharded Data Parallel (FSDP), Sequence Parallelism (SP), and Expert Parallelism (EP)) are designed to be fully composable and can be flexibly applied to different components of an omni-modal LLM. This composability is particularly advantageous for multimodal training, where different modalities or architectural components may benefit from distinct parallelism techniques. For example, vision encoders can adopt FSDP or standard data parallelism depending on their scale, while language backbones can leverage a combination of EP for MoE layers and SP to support long-context processing. This fine-grained flexibility enables efficient and scalable training across diverse model architectures.
Figure 3 provides an overview of VeOmni’s n-D parallelism design. The left side of the figure illustrates the communication and computation flow for a unified multimodal model, where different modality-specific encoders process their respective inputs and distribute intermediate features to downstream modules. The right side visualizes the application of hybrid parallelism strategies across different parts of the model, demonstrating how various parallelism techniques coexist within a single training configuration. To support this flexible parallelism composition, VeOmni introduces a unified abstraction for distributed training based on a global device mesh. Unlike approaches [51, 41] that require manually managing multiple process groups, VeOmni significantly simplifies the configuration and coordination of n-D parallelism(Appendix B.4). This not only reduces the complexity of process group management but also improves extensibility, making it easier to adapt and extend to future parallelism strategies or new model components.
@@ 3.2.5 Other System Optimization Strategies
In addition to the aforementioned parallelism strategies, VeOmni also incorporates a wide range of other system-level optimizations, following the key design principle of VeOmni, which is the decoupling of these optimization implementations from the model’s core logic. This modular architecture allows for seamless integration, enabling these system-level enhancements to be readily applied across various models with minimal to no modification of the model code. These optimizations include, but are not limited to, the following:
@@
  • Efficient Kernels. To maximize training throughput, VeOmni incorporates a suite of highly optimized operator kernels, including RMSNorm, LayerNorm, RoPE, SwiGLU, and in-place CrossEntropy from Liger-kernel [23], along with FlashAttention [9, 8, 49] and MoE-specific operations [4, 74]. These kernels are carefully engineered for both high performance and broad compatibility, enabling efficient execution across diverse transformer-based architectures and model variants.
@@
  • Efficient Distributed Checkpointing. VeOmni leverages ByteCheckpoint [63] to enable efficient checkpoint saving and resumption across varying distributed configurations with minimal overhead. Beyond facilitating elastic training and enhancing cluster utilization, our framework further extends ByteCheckpoint to support multimodal models. This extension ensures consistent and reliable saving and loading of heterogeneous model components.
  • Meta Device Initialization. VeOmni supports model initialization on the meta device for large models without allocating physical memory during the definition phase. After the model structure is instantiated on the meta device, we perform parameter sharding and parallel loading by converting parameters into the DTensor format, significantly accelerating the initialization and loading processes for large-scale models.

### v3

@@ 1 Introduction
Extending existing training frameworks to omni-modal LLMs is non-trivial. To address the growing parameter sizes, modern training frameworks leverage various distributed training strategies [52, 40, 32, 41] to alleviate computation and memory bottlenecks. For example, Megatron-LM [52, 41] provides optimized transformer blocks [62] with advanced parallelization strategies like tensor parallelism (TP) and pipeline parallelism (PP). However, as omni-modal architectures become complex in both functionality and parameter size, directly applying these techniques to omni-modal LLMs often leads to load imbalance and poor scalability [17], due to the current frameworks tightly couple model definition with parallel logic. This entanglement hiders generalization to more diverse model architectures. Although recent efforts [24, 79, 17] attempt to mitigate the inefficiencies of omni-modal training, they still suffer from the coupling of communication and computation, resulting in significant engineering cost and limited extensibility.
To efficiently train large omni-modal LLMs, VeOmni introduces model-centric distributed recipes that decouple model definition from parallel logic. As shown in Figure 1, distributed strategies such as fully sharded data parallel (FSDP), hybrid sharded data parallel (HSDP) [80, 36], sequence parallelism (SP) [29], and expert parallelism (EP) [76] can be applied to model blocks via a high-level parallel plan API. This design effectively separates communication from computation and has been validated in recent LLM training systems [36]. VeOmni supports flexible composition of parallel strategies (e.g., FSDP+SP for 2D and FSDP+SP+EP for 3D parallelism), enabling a range of training recipes tailored to dense or MoE models. By decoupling parallel logic, new modality-specific modules can be integrated with minimal engineering effort, as computation modules need not account for distributed concerns. In addition, VeOmni provides a light-weight interface for customizing omni-modal LLMs, allowing users to easily modify the modality-specific encoders and decoders.
@@
  1. 1.
    We propose model-centric distributed recipes that efficiently scale omni-modal training using composable n-D parallelism with minimal engineering overhead.
  2. 2.
    We introduce a light-weight configuration interface for flexibly customization of omni-modal LLMs, supporting both multi-modal understanding and generation.
  3. 3.
    We demonstrate VeOmni’s competitive efficiency and scalability across 8–128 GPUs on models ranging from 7B to 72B parameters under omni-modal scenarios.
@@ 2.1 Multi-Modal and Omni-Modal LLMs
Recent advances in large language models (LLMs) have increasingly focused on developing multi-modal and omni-modal capabilities. The prevailing approaches include incorporating modality-specific encoders into the input space of the LLMs for multi-modal understanding [39, 66, 74, 37, 19], and attaching generative decoders to the output space for multi-modal generation [27, 60] or embodied action prediction [2, 31]. More recently, omni-modal LLMs that support unified understanding and generation across modalities have emerged, aiming to align arbitrary modality features with language in a shared latent space. Based on a unified paradigm of auto-regressive modeling over multi-modal embeddings, these models differ significantly in how the modality features are encoded and decoded. Some [57, 68] adopt discrete-token generation pipelines based on VQ-VAE series [16, 47]. Some [65, 25, 55] employ continuous-token generation via latent diffusion models [49, 44]. Others [12, 73, 82] explore hybrid architectures that combine diffusion-based generation with auto-regressive decoding. In addition, some works proposed alternative next-token prediction schemes, such as masked generation [34], next patch prediction [42], next scale prediction [59], next block prediction [48], offering increased flexibility beyond standard auto-regressive decoding. Given this diverse landscape of model architectures, VeOmni offers a flexible and extensible framework for building and scaling new modeling paradigms.
@@ 2.2 LLM Training Frameworks
The landscape of large language model training frameworks has evolved significantly to address the computational demands of scaling. For pure text training, several mature frameworks have established themselves as industry standards. Megatron-LM [52, 41] pioneered optimized transformer blocks with advanced parallelization strategies including tensor parallelism (TP) and pipeline parallelism (PP), becoming the foundation for many subsequent frameworks. Colossal-AI [33] extends this paradigm by offering comprehensive 3D parallel training strategies that combine data, tensor, and pipeline parallelism for enhanced scalability. NeMo [51] provides an end-to-end cloud solution specifically designed for training large-scale LLMs with enterprise-grade features, while TorchTitan [36] and veScale [63] represent the newer generation of PyTorch-native frameworks that emphasize auto-parallelization and simplified programming models.
@@
3 VeOmni: Scalable Omni-Modal Training Framework
@@
3.1 Light-Weight Customization for Omni-Modal LLMs
Figure 2: Composite architecture for omni-modal LLMs. The architecture consists of three fully decoupled modules: encoder, foundation model, and decoder.
To support training across arbitrary modalities, we propose a plug-and-play architecture that allows any combination of multimodal encoders and decoders to be flexibly attached to the input and output sides of a foundation model, as shown in Figure 2. This modular and fully decoupled design decouples system-level operations from model-specific computation, allowing the streaming pipeline to operate independently of the model pipeline. Therefore, VeOmni enables easy extension and scaling of diverse omni-modal LLMs.
@@ 3.1 Light-Weight Customization for Omni-Modal LLMs
In the training phase, for input modalities, each encoder implements an lm_encode function that encodes raw modality data into token embeddings, which are then inserted into the input embeddings of the foundation model. Similarly, decoder modules implement the lm_encode function to provide training-time inputs representing the target output tokens. The final outputs from the foundation model are decoded into the target modality using lm_head function, completing the end-to-end mapping.
In the inference phase, for each prediction step, the output hidden states is passed through the corresponding modality decoder’s lm_embed function, which generates embeddings that are used as inputs for the foundation model’s next token prediction. Once all tokens have been predicted, the decoder’s lm_generate function is invoked to generate the final modality-specific output. The details of the training and inference processes are listed in Appendix B.3.
@@ 3.2 Model-Centric Distributed Recipe Zoo
To support large-scale distributed training of omni-modal LLMs, VeOmni offers a modular suite of distributed training strategies, including fully sharded data parallel (FSDP), sequence parallelism (SP), and expert parallelism (EP), designed to handle challenges such as training with ultra-long sequences and efficiently scaling large mixture-of-experts (MoE) models. VeOmni enables users to freely compose distributed training recipes tailored to their model architectures and modality-specific requirements, without needing to manage low-level implementation details. The framework is designed with the following principles:
@@
    Non-intrusive distributed training APIs. VeOmni decouples the model architecture from the underlying distributed training mechanisms, enabling users to compose omni-modal models flexibly without modifying low-level parallelization code.
@@
    Support for long sequence training. VeOmni accommodates training workloads involving extremely long sequences across different modalities, which is a common challenge in omni-modal training.
@@
3.2.1 Fully Sharded Data Parallel for Large Model Training
@@
Fully sharded data parallel (FSDP) is a highly efficient implementation of the zero redundancy optimizer (ZeRO) [46] in PyTorch. Its primary advantage is the significant reduction in memory required on each GPU during training. By sharding a model’s parameters, gradients, and optimizer states across all available devices, FSDP allows for the training of extremely large models that would otherwise exceed the memory capacity of a single GPU. A key advantage of FSDP is its non-intrusive design, which decouples the model’s architecture from the parallelization strategy. This means developers can focus on designing their models without needing to modify the underlying code to accommodate the distributed training setup. This non-intrusive nature makes FSDP exceptionally well-suited for training omni-modal LLMs. Since these models are defined by their complex and heterogeneous architectures, FSDP’s ability to parallelize training without requiring any changes to the model’s code makes it an ideal solution. VeOmni integrates both FSDP1 [80] and FSDP2 [36] as fundamental components of its distributed training stack, and provides a unified API for easy configuration. More details in Appendix B.5.
@@
To further improve training efficiency, VeOmni integrates hybrid sharded data parallel (HSDP), an extension of FSDP designed to minimize communication overhead. HSDP utilizes a 2D device mesh, combining FSDP within “shard groups” and distributed data darallel (DDP) across “replicate groups”. This hybrid approach drastically cuts down on inter-node communication, enabling even greater scalability. The switch to the more efficient HSDP is as simple as changing one parameter in the configuration, with no modifications to the underlying code required.
3.2.2 Sequence Parallelism for Long Sequence Training
@@
With the advancement of state-of-the-art omni-modal LLMs, the sequence lengths required for training have grown significantly, particularly in domains such as high-resolution image or video understanding and generation [66, 78, 18]. This trend toward longer sequences poses substantial challenges in terms of both computational cost and memory consumption. Efficiently addressing the demands of long sequence training is therefore critical for scaling model capacity and unleashing the full potential of omni-modal architectures. To address this, VeOmni adopts DeepSpeed-Ulysses [29], a highly efficient sequence parallelism technique specifically designed for transformer training on ultra-long sequences. DeepSpeed-Ulysses splits activations along the sequence dimension and strategically orchestrates all-to-all communications during attention computation, ensuring that communication volume remains constant when both sequence length and device count are scaled proportionally. This design enables high efficiency and scalability for long-sequence training. A core advantage of VeOmni is its provision of a non-intrusive interface for deploying DeepSpeed-Ulysses. Developers can seamlessly leverage DeepSpeed-Ulysses without needing to modify model-building code or directly interact with the underlying implementation details. In particular, VeOmni enhances the implementation of FlashAttention [9] by integrating DeepSpeed-Ulysses while keeping fully compatible with the default attention, enabling the straightforward adoption of advanced attention acceleration methods within sequence parallel workflows. More details in Appendix B.6.
@@ 3.2.3 Expert Parallelism for MoE Model Scaling
3.2.4 Composable n-D Parallel Training
Figure 3: Overview of VeOmni’s n-D parallelism. The figure on the left illustrates the data flow of VeOmni during the training of omni-modal LLMs. The encoder for each modality processes its respective inputs and then scatters the features to the corresponding ranks through an all-to-all communication operation. The right figure shows the 3D parallelism of VeOmni, where attention blocks apply HSDP with a data sharded size of 2 and a data replicate size of 2, inputs apply sequence parallelism with a size of 2, and mixture-of-experts blocks apply expert parallelism and FSDP with an EP size of 2 and a data sharded size of 2.
In VeOmni, core parallelism strategies (i.e., fully sharded data parallel (FSDP), sequence parallelism (SP), and expert parallelism (EP)) are designed to be fully composable and can be flexibly applied to different components of an omni-modal LLM. This composability is particularly advantageous for omni-modal training, where different modalities or architectural components may benefit from distinct parallelism techniques. For example, vision encoders can adopt FSDP or standard data parallelism depending on their scale, while language backbones can leverage a combination of EP for MoE layers and SP to support long sequence processing. This fine-grained flexibility enables efficient and scalable training across diverse model architectures.
Figure 3 provides an overview of VeOmni’s n-D parallelism design. The left side of the figure illustrates the communication and computation flow for a omni-modal model, where different modality-specific encoders process their respective inputs and distribute intermediate features to downstream modules. The right side visualizes the application of n-D parallelism strategies across different parts of the model, demonstrating how various parallelism techniques coexist within a single training configuration. To support this flexible parallelism composition, VeOmni introduces a unified abstraction for distributed training based on a global device mesh [36]. Unlike approaches [52, 41] that require manually managing multiple process groups, VeOmni significantly simplifies the configuration and coordination of n-D parallelism (Appendix B.4). This not only reduces the complexity of process group management but also improves extensibility, making it easier to adapt and extend to future parallelism strategies or new model components.
@@ 3.2.5 Other System Optimization Strategies
In addition to the aforementioned parallelism strategies, VeOmni also incorporates a wide range of other system-level optimizations, following the key design principle of VeOmni, which is the decoupling of these optimization implementations from the model’s computation logic. This modular architecture allows for seamless integration, enabling these system-level enhancements to be readily applied across various model architectures with minimal to no modification of the model code. These optimizations include, but are not limited to, the following strategies:
@@
  • Efficient Kernels. To maximize training throughput, VeOmni incorporates a suite of highly optimized operator kernels, including RMSNorm, LayerNorm, RoPE, SwiGLU, and in-place cross-entropy from liger-kernel [23], along with FlashAttention [9, 8, 50] and MoE-specific operations [4, 76]. These kernels are carefully engineered for both high performance and broad compatibility, enabling efficient execution across diverse transformer-based architectures and model variants.
@@
  • Efficient Distributed Checkpointing. VeOmni leverages ByteCheckpoint [64] to enable efficient checkpoint saving and resumption across varying distributed configurations with minimal overhead. Beyond facilitating elastic training and enhancing cluster utilization, our framework further extends ByteCheckpoint to support omni-modal models. This extension ensures consistent and reliable saving and loading of heterogeneous model components.
  • Meta Device Initialization. VeOmni supports model initialization on the meta device for large models without allocating physical memory during initialization. After instantiating the model on the meta device, we perform parameter sharding and parallel loading by converting parameters into the DTensor format [36], significantly accelerating the initialization and loading processes for large-scale models.